OpenStudy (anonymous):

Y: 2x(3+y)=12 i don't know how to solve this one

OpenStudy (anonymous):

start by distributing

OpenStudy (anonymous):

then what?

OpenStudy (anonymous):

what is your answer after you distribute?

OpenStudy (anonymous):

6x+2xy=12

OpenStudy (anonymous):

CAN'T YOU DIVIDE BY 2X?

OpenStudy (anonymous):

next i would get xy by itself

OpenStudy (anonymous):

THEN SUBTRACT 3 ON BOTH SIDES

OpenStudy (anonymous):

i have xy=1/2 not sure what to do now

OpenStudy (anonymous):

how did you get rid of the other x?

OpenStudy (anonymous):

2x divide 2 by 2

OpenStudy (anonymous):

nope would still be x

OpenStudy (anonymous):

ok i should be good from here thnx

OpenStudy (anonymous):

2xy=12-6x xy=6-3x y=?

OpenStudy (anonymous):

I would have divided by 2x first. Saves steps later and simplifies the problem.

OpenStudy (anonymous):

either way still gets to the same spot

OpenStudy (anonymous):

2x(3+y)=12 3 + y = 6/x y = 6/x - 3 done and done

OpenStudy (anonymous):

except this way is faster and easier. ;)

OpenStudy (anonymous):

thats what i did!

OpenStudy (anonymous):

thanks all of you :)

OpenStudy (anonymous):

fair enough